# Poincare field theory for massless particles

### v2

@@ Abstract:
Our main proposition is that field equations for all spins can be obtained from Casimir eigenvalue equations for Poincare group. We have already confirm that statement for massive scalar, spinor and vector fields in Ref.[1]. In the present article we are going to confirm this statement for massless vector, second and fourth rank tensor fields. In particular we will obtain Maxwell equations and Einstein equation in weak field approximation.
As is well known, Wigner defined particles as irreducible representation of Poincare group [2, 3]. But, as Weinberg noted in [4] irreducible representations for massless vector field with helicities ±$\pm 1$ do not exist. In the present article we will conform this statement for a wide class of massless fields. They are not Lorentz invariant since their Lorentz transformations have additional term in the form of gauge transformations. Such fields can appear in the theory only in the form which do not depend on corresponding gauge parameters. These forms are our equations of motion and they are by definition gauge invariant. So, the massless case is significantly different from massive one. In the end we will show that our approach can reproduce main contributions from well known articles with highest helicity.
@@ 1 Introduction
In the previous article [1] we introduced equations of motion for massive particles with arbitrary spin as eigenvalue equations for Casimir operators of Poincare group. Here we will develop appropriate approach for massless particles. Although in massless case there is one Casimir operator (helicity), covariant form of its definition contains Pauli-Lubanski vector $W_{a}$ which has three independent components (since $P^{a}W_{a}=0$). So, we have three equations and their compatibility produces significant differences from massive case.
As in the massive case the fields are functions which transforms with some representation of Poincare group. For arbitrary representations we will introduce notation $\Psi^{A}(x)$ where $A$ contains the set of corresponding indices.
@@
We will work in the frame of standard momentum which remarkably simplifies all considerations. First we will solve eigenvalue problem for component of spin operator $(S_{12})^{A}{}_{B}=\sigma(W_{0})^{A}{}_{B}=(W_{3})^{A}{}_{B}$. Corresponding eigenvalues we can interpret as helicities $\lambda_{i}$ and eigenvectors $\Psi^{A}_{i}$ as potential irreducible representations of little group. But in some important cases components of Pauli-Lubanski vector do not commute. So, there is no common eigenvectors for all components. Even more, components $W_{1}$ and $W_{2}$ in the frame of standard momentum do not annihilate eigenvectors $\Psi^{A}_{i}$, which means that $\Psi^{A}_{i}$ are not representations of Poincare group.
@@
This feature has unexpected consequences discovered by Weinberg [4] that Lorentz transformations of some physically relevant fields (like vector and symmetric second ranktensor fields) have additional term in the form of gauge transformation. So, for these fields we can obtain actions as gauge invariant expressions of eigenvectors $\Psi^{A}_{i}$.
Therefore, gauge dependent fields are not Lorentz invariant and so they are not representation of Poincare group. But corresponding field strength and Lagrangian (gauge invariant expression of gauge fields) are representations of Poincare group. So, for gauge fields we will introduce name incomplete representation of Poincare group (up to gauge transformations) and for corresponding irreducible part incomplete irreducible representation of Poincare group.
@@
As a consequences, Wigner’s definition of particles [2, 3] as irreducible representations of Poincare group, which valid in massive case, in massless case must have an improvement. We can define physical massless fields as a class of equivalence, where two fields are equivalent if they relate by gauge transformation. The action is expression which depend on class of equivalence.
The main idea of this work is to show that principle field equations produce all known field equations. So, we will first obtain standard equations for helicities $1$ and $2$. There is also significant interest in massless fields with higher spin. The theory of particles with highest spin was developed by Fierz and Pauli [5]. In Ref. [2] and [3] it has been shown that the one-particle states is an irreducible, unitary representation of the Poincare group. The set of a minimum auxiliary traceless, symmetric tensor fields was proposed in Ref.[6]. Since it is convenient to have an action for a field theory in order to perform quantization, equations of motion and action was found in Ref.[7]. There is also interesting approach based on a hierarchy of generalized Christoffel symbols by de Wit and Freedman [8]. In all of the above articles, fields under consideration satisfy Fierz-Pauli constraints [5] which are the basis of these articles. It means that field is totally symmetric tensor of rank $n$, which is traceless, divergence free and satisfying the Klein-Gordon equation.
In the present article we offer more fundamental approach. For arbitrary massless tensor we developed procedure to find all possible helicities and structure od tensor for each given helicity. As a particular case, for given tensor of rank $n$ highest helicity is totally symmetric, traceless, divergence free and satisfying the Klein-Gordon equation. So, our highest helicity satisfy Fierz-Pauli constraints [5]. We also considered all non-highest helicity and offer theirs physical interpretation, see sections 7. and 8.
@@ 2 Principle field equations
In this section we will introduce principle field equations. They are defined only on the basis of symmetry group of the theory and they are formulated with the help of Casimir invariants. In this article we start with massless Poincare group and introduce principle field equations which valid for arbitrary helicity.
@@ 2.3 Principle field equations
With the help of expression (1) and representation of momenta and Pauli-Lubanski vector we can postulate principle field equations for massless fields $\Psi^{A}(x)$ with helicity $\lambda$
@@
We claim that all massless free field equations with arbitrary helicity, considered in classical field theory [4, 9, 10, 11, 12] can be obtained from this equation. In the present article we will confirm this statement for most important massless fields.
@@ 4.1 Eigenvalues and projection operators
Note that eigenspace $\Xi_{0}$ corresponding to $\lambda_{0}=0$ is two dimensional $d_{0}=dim\Xi_{0}=2$ and consequently corresponding eigenvector carries two degrees of freedom. The eigenspaces $\Xi_{\pm 1}$ corresponding to $\lambda_{\pm 1}=1$ are one dimensional $d_{\pm 1}=dim\Xi_{\pm 1}=1$ and consequently each of corresponding eigenvectors carries one degree of freedom.
@@ 6.2 General relativity in weak field approximation
In order to get field equation we need gauge invariant function of tensor $h^{ab}(x)$. There are several solutions: Riemann, Ricci and scalar curvatures as well as Einstein tensor in weak field approximation. They can be obtained as small perturbation around flat metric using relation $g_{ab}=\eta_{ab}+h_{ab}$, see Refs.[11, 12].
Since equation of motion with respect to $h^{ab}$ has two indices we can chose Ricci or Einstein tensor. We are going later to include interaction and so we will take Einstein tensor for free field equation (for expression for Einstein tensor see Sec.8)
$G^{ab}(x)=R^{ab}-\frac{1}{2}\eta^{ab}R=\partial_{c}\partial^{b}h^{ac}-\partial^{2}h^{ab}-\partial^{a}\partial^{b}h+\partial^{a}\partial_{c}h^{cb}-\eta^{ab}\Big{(}\partial_{c}\partial_{d}h^{cd}-\partial^{2}h\Big{)}=0\,.\qquad$ (9)
Similarly as in the case of vector field it can be shown (see for example Ref.[12]) that field $h^{ab}$ which satisfy this equation has two degrees of freedom. The symmetric tensor $h^{ab}$ has $10$ components. To fix gauge invariance (8) we can use $4$ gauge fixings conditions $\partial_{a}h^{ab}=\frac{1}{2}\partial^{b}h^{a}{}_{a}$ which kills $4$ component of tensor field and after that equation of motion takes the form $\partial^{2}\Big{(}h^{ab}-\frac{1}{2}\eta^{ab}h\Big{)}=0$. But now both gauge fixing and field equation are invariant under additional special gauge transformation $\delta h^{ab}(x)=\partial^{a}\Omega_{0}^{b}(x)+\partial^{b}\Omega_{0}^{a}(x)$, such that $\partial^{2}\Omega_{0}^{a}=0$. Corresponding gauge fixing kills additional $4$ degree of freedom and consequently gravitation field $h^{ab}(x)$ has two degrees of freedom.
Field equation (9) can be obtained from the action
@@ 7.2.1 Symmetric tensor with helicities $\lambda=\pm 1$
$\delta\varphi^{ab}(x)(x)=\partial^{a}\partial^{b}\Big{(}\alpha\omega_{+}(x)+\beta\omega_{-}(x)\Big{)}=\partial^{a}\partial^{b}\omega(x)\,.\qquad$ (15)
@@ 7.2.2 Antisymmetric tensor with helicities $\lambda=\pm 1$ and Maxwell equations
As symmetric second rank tensor with helicities $\lambda=\pm 1$ describe electromagnetic interaction through field strength, we can expect that fourth rank tensor with helicities $\lambda=\pm 2$ and appropriate symmetries describe gravitation interaction through Riemann tensor. In the case of fourth rank tensor we have $A\to(abcd)$ and $\Psi^{A}\to T^{abcd}$. In this section to simplify notation we will keep index $A$ instead of four indices $abcd$. When we include symmetry properties we will go back to the notation with indices $abcd$.
@@ 8.1 Fourth rank tensor
Here $(P_{i})^{A}{}_{B}$ is sum of all terms $\pi_{i_{1}}\pi_{i_{2}}\pi_{i_{3}}\pi_{i_{4}}$ such that $\sum_{m=1}^{4}i_{m}=i$. We count $\pi_{+}$ as $\pi_{1}$ and $\pi_{-}$ as $\pi_{-1}$. Let us stress that order of multipliers is important, so that for example $\pi_{+}\pi_{0}\pi_{+}\pi_{0}$ is different from $\pi_{0}\pi_{+}\pi_{+}\pi_{0}$. Note that as a consequences of these rules, and taking into account that $\pi_{0},\pi_{+}$ and $\pi_{-}$ are projectors, all expressions $P_{i}$ where $i=0,\pm 1,\pm 2,\pm 3,\pm 4$ are projectors also, $P_{i}P_{j}=\delta_{ij}P_{i}$.
@@ 9 Massless tensors with arbitrary rank
Here $(P_{k})^{A}{}_{B}$ is sum of all terms $\pi_{i_{1}}\pi_{i_{2}}\cdots\pi_{i_{n}}$ such that $\sum_{m=1}^{n}i_{m}=k$ where we count $\pi_{+}$ as $\pi_{1}$ and $\pi_{-}$ as $\pi_{-1}$. Note that since $\pi_{0}$, $\pi_{+}$ and $\pi_{-}$ are projectors all expressions $P_{k}$ are projectors also, because $P_{k}P_{q}=\delta_{kq}P_{k}$ where $(k,q=0,\pm 1,\pm 2,\cdots,\pm n)$. The multiplication factor $k$ in (5) is consequence of combinatorics.
@@ 9.3 Dimension of subspaces $d_{k}$
Dimension of subspace after projection with $\pi_{0}$ is $2$ while dimensions of subspaces after projections with $\pi_{+}$ and $\pi_{-}$ are $1$. Since $d_{k}=d_{-k}$ we have following expressions for $d_{k}$.
@@ 10 Components with highest helicity
In previous sections we showed that our principle equation reproduces all examples with lower helicities $\lambda=0,1,2$. It is interesting to compare our approach to the known literature with higher helicity examples $\lambda\geq 3$,
The general description of free fields with helicity $n$ has been developed first by Fronsdal [7]. We are going to show that this is particular case of our approach for highest helicity.
@@ 10.2 Highest helicity tensor $h_{A}(k)$ symmetric under space inversion in the frame of standard momentum
11 Relation with literature
In this section, using our approach, we will derive some basic results of well known articles for highest helicities.
11.1 Conditions for highest helicity in coordinate representation satisfy Fierz-Pauli constraints
Let us go from standard momentum frame to arbitrary frame and then to coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$. Then we will consider field $h^{A}(x)$ in coordinate representation

### v3

@@ Abstract:
Our main proposition is that the field equations for all spins can be derived from the Casimir eigenvalue equations of the Poincaré group. We have already confirmed this claim for massive scalar, spinor, and vector fields in Ref.[1]. In this article, we extend this analysis to massless fields—specifically, the vector field, as well as second and fourth rank tensor fields and finally for arbitrary rank tensor fields. In particular, we will derive the Maxwell equations, the Einstein equations in the weak-field approximation as well as Fronsdal equation.
As is well known, defined particles as irreducible representations of the Poincaré group [2, 3]. However, as Weinberg noted in [4], irreducible representations for massless vector fields with helicities ±$\pm 1$ do not exist. In this article, we confirm this statement for a broad class of massless fields. Unlike their massive counterparts, these fields are not Lorentz invariant because their Lorentz transformations include an additional gauge transformation term. Consequently, such fields can only appear in the theory in forms that are independent of the corresponding gauge parameters. These gauge invariant forms constitute our equations of motion. Thus, the massless case differs fundamentally from the massive one.
Finally, we demonstrate that our approach not only reproduces key results from well-known studies on highest helicity fields but also extends to other helicities.
@@ 1 Introduction
In our previous work [1], we derived the equations of motion for massive particles of arbitrary spin by formulating them as eigenvalue equations for the Casimir operators of the Poincaré group. In this paper, we generalize this framework to massless particles, where new subtleties arise due to gauge invariance.
While only one Casimir operator (helicity) exists for massless particles, its covariant definition involves the Pauli-Lubanski vector $W_{a}$, which possesses three independent components (due to the constraint $P^{a}W_{a}=0$). Thus, we obtain three equations, and their compatibility conditions lead to significant distinctions from the massive case.
As in the massive scenario, the fields are functions transforming under a representation of the Poincaré group. For a general representation, we denote these fields as $\Psi^{A}(x)$, where $A$ encompasses the relevant set of indices.
@@
We will work in the standard momentum frame, which significantly simplifies our analysis. First, we solve the eigenvalue problem for the spin operator component $(S_{12})^{A}{}_{B}=\sigma(W_{0})^{A}{}_{B}=(W_{3})^{A}{}_{B}$. The resulting eigenvalues can be interpreted as helicities $\lambda_{i}$ and the corresponding eigenvectors $\Psi^{A}_{i}$ serve as candidates for irreducible representations. However, in certain important cases, the components of the Pauli-Lubanski vector fail to commute. Consequently, no common eigenvectors exist for all components. Moreover, in the standard momentum frame, the operators $W_{1}$ and $W_{2}$ do not annihilate eigenvectors $\Psi^{A}_{i}$, implying that $\Psi^{A}_{i}$ cannot represent genuine irreducible representations of the Poincaré group.
@@
This feature has unexpected consequences, as discovered by Weinberg [4], where the Lorentz transformations of certain physically relevant fields (such as vector and symmetric second-rank tensor fields) acquire an additional term in the form of a gauge transformation. Consequently, for these fields, we can construct actions as gauge-invariant expressions involving the eigenvectors $\Psi^{A}_{i}$.
Thus, gauge-dependent fields are not Lorentz invariant and therefore do not constitute proper representations of the Poincaré group. However, the corresponding field strengths and Lagrangians—being gauge-invariant expressions of the gauge fields—do form representations of the Poincaré group. For this reason, we introduce the term incomplete representation of the Poincaré group (up to gauge transformations) for gauge fields, and incomplete irreducible representation of the Poincaré group for their corresponding irreducible components.
@@
As a consequence, Wigner’s definition of particles [2, 3] as irreducible representations of the Poincaré group, while valid for massive cases, requires refinement in the massless case. Instead, we can define physical massless fields as equivalence classes, where two fields are considered equivalent if they are related by a gauge transformation. The action is then an expression that depends only on the equivalence class.
The main objective of this work is to demonstrate that principle field equations give rise to all known field equations. To this end, we first derive the standard equations for helicities $1$ and $2$. There is also considerable interest in massless fields with higher spin. The theory of particles with highest spin was originally developed by Fierz and Pauli [5]. In Ref. [2] and [3] , it was shown that the one-particle states correspond to irreducible, unitary representations of the Poincaré group.
A minimal set of auxiliary traceless, symmetric tensor fields was later proposed in Ref.[6]. Since having an action is essential for field quantization, the equations of motion and corresponding action were derived in Ref.[7]. Another notable approach, based on a hierarchy of generalized Christoffel symbols, was introduced by de Wit and Freedman [8]. All the above studies rely on the Fierz-Pauli constraints [5] which serve as their foundation. These constraints require that the field be a fully symmetric tensor of rank $n$, while also being traceless, divergence-free, and satisfying the Klein-Gordon equation.
In this work, we present a more fundamental approach to analyzing massless tensor fields. We develop a systematic procedure to determine all possible helicity states and the corresponding tensor structures for any given helicity.
As a special case, for a rank $n$ tensor field, the highest helicity state corresponds to a fully symmetric, traceless, and divergence-free tensor that satisfies the Klein-Gordon equation. Thus, our highest helicity states naturally satisfy the Fierz-Pauli constraints [5]. Beyond the highest helicity, we also investigate non-maximal helicity states and provide their physical interpretation (see Sections 7 and 8 and 12).
@@ 2 Principle field equations
In this section, we introduce the fundamental field equations of the theory. These equations are constructed solely from the symmetry group properties and are formulated using the Casimir invariants.
In the present work, we begin with the massless Poincaré group and derive fundamental field equations that hold for arbitrary helicity states.
@@ 2.3 Principle field equations
Using the expression (1) along with the momentum and Pauli-Lubanski vector we postulate principle field equations for massless fields $\Psi^{A}(x)$ with helicity $\lambda$
@@
We demonstrate that all massless free field equations with arbitrary helicity in classical field theory [4, 9, 10, 11, 12] can be derived from this fundamental equation. In this work, we explicitly verify this claim for the most physically significant massless fields.
@@ 4.1 Eigenvalues and projection operators
We note that eigenspace $\Xi_{0}$ corresponding to the eigenvalue $\lambda_{0}=0$ is two dimensional $d_{0}=dim\Xi_{0}=2$ implying that its associated eigenvector carries two degrees of freedom. Similarly, the eigenspaces $\Xi_{\pm 1}$ corresponding to eigenvalues $\lambda_{\pm 1}=1$ respectively are both one dimensional $d_{\pm 1}=dim\Xi_{\pm 1}=1$. Consequently, each of their associated eigenvectors carries a single degree of freedom.
@@ 6.2 General relativity in weak field approximation
In order to get field equation we need gauge invariant function of tensor $h^{ab}(x)$. It is linear combination of massless Klein-Gordon equation and supplementary conditions $G^{b}$
$\partial^{2}h_{ab}=0\,,\qquad G^{b}=\partial_{a}h^{ab}-\frac{1}{2}\partial^{b}h^{a}{}_{a}=0\,.$ (9)
The supplementary conditions are in fact gauge fixing which kill all unwanted degrees of freedom.
These equations can be combined into one equation as linear combination of Klein-Gordon equation and supplementary conditions
$\partial^{2}h_{ab}+\alpha\Big{(}\partial_{a}G_{b}+\partial_{b}G_{a}\Big{)}=0\,.$ (10)
It must be gauge invariant and using gauge transformation of gauge fixing condition $\delta G_{a}=\partial^{2}\Omega_{a}$ we obtain $\alpha=-1$.
Consequently, equation of motion takes a form of Ricci curvature
$R_{ab}=\partial^{2}h_{ab}-\Big{(}\partial_{a}G_{b}+\partial_{b}G_{a}\Big{)}=\partial^{2}h_{ab}-\partial_{a}\partial^{c}h_{cb}-\partial_{b}\partial^{c}h_{ca}+\partial_{a}\partial_{b}h^{c}{}_{c}=0\,.$ (11)
Then scalar curvature is
$R=\eta^{a}R_{ab}=\partial^{2}h-2\partial^{a}G_{a}=2\Big{(}\partial^{2}h-\partial^{a}\partial^{b}h_{ab}\Big{)}\,,$ (12)
and Einstein tensor for free field equation
$G^{ab}(x)=R^{ab}-\frac{1}{2}\eta^{ab}R=\partial^{2}\Big{(}h_{ab}-\frac{1}{2}\eta^{ab}h\Big{)}-\Big{(}\partial_{a}G_{b}+\partial_{b}G_{a}\Big{)}+\eta^{ab}\partial^{c}G_{c}$
$=\partial_{c}\partial^{b}h^{ac}-\partial^{2}h^{ab}-\partial^{a}\partial^{b}h+\partial^{a}\partial_{c}h^{cb}-\eta^{ab}\Big{(}\partial_{c}\partial_{d}h^{cd}-\partial^{2}h\Big{)}=0\,.\qquad$ (13)
All these tensors can be obtained as small perturbation around flat metric using relation $g_{ab}=\eta_{ab}+h_{ab}$, see Refs.[11, 12].
Field equation (6.2) can be obtained from the action
@@
Similarly as in the case of vector field it can be shown (see for example Ref.[12]) that field $h^{ab}$ which satisfy this equation has two degrees of freedom. The symmetric tensor $h^{ab}$ has $10$ components. To fix gauge invariance (8) we can use $4$ gauge fixings conditions $G^{b}=\partial_{a}h^{ab}-\frac{1}{2}\partial^{b}h^{a}{}_{a}=0$ which kills $4$ component of tensor field and after that equation of motion takes the form $\partial^{2}\Big{(}h^{ab}-\frac{1}{2}\eta^{ab}h\Big{)}=0$. But now both gauge fixing and field equation are invariant under additional special gauge transformation $\delta h^{ab}(x)=\partial^{a}\Omega_{0}^{b}(x)+\partial^{b}\Omega_{0}^{a}(x)$, such that $\partial^{2}\Omega_{0}^{a}=0$. Corresponding gauge fixing kills additional $4$ degree of freedom and consequently gravitation field $h^{ab}(x)$ has two degrees of freedom.
@@ 7.2.1 Symmetric tensor with helicities $\lambda=\pm 1$
$\delta\varphi^{ab}(x)=\partial^{a}\partial^{b}\Big{(}\alpha\omega_{+}(x)+\beta\omega_{-}(x)\Big{)}=\partial^{a}\partial^{b}\omega(x)\,.\qquad$ (15)
@@ 7.2.2 Antisymmetric tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Since a symmetric second-rank tensor field with helicities $\lambda=\pm 1$ describes electromagnetic interactions through the field strength tensor, we naturally expect that a fourth-rank tensor field with helicities $\lambda=\pm 2$ and appropriate symmetry properties should describe gravitational interactions through the Riemann curvature tensor.
For the fourth-rank tensor case, the field index $A$ corresponds to a quadruple index $A\to(abcd)$ with $\Psi^{A}\to T^{abcd}$. In this section, we will maintain the compact notation using index $A$ for simplicity, reverting to the explicit index notation $abcd$ only when discussing specific symmetry properties.
@@ 8.1 Fourth rank tensor
Here $(P_{i})^{A}{}_{B}$ is defined as the sum of all possible terms $\pi_{i_{1}}\pi_{i_{2}}\pi_{i_{3}}\pi_{i_{4}}$ such that $\sum_{m=1}^{4}i_{m}=i$ where we assign $\pi_{+}\equiv\pi_{1}$ and $\pi_{-}\equiv\pi_{-1}$. The sequence of multipliers is significant. For example $\pi_{+}\pi_{0}\pi_{+}\pi_{0}$ is different from $\pi_{0}\pi_{+}\pi_{+}\pi_{0}$. Since $\pi_{0},\pi_{+}$ and $\pi_{-}$ are projectors, it follows that all $P_{i}$ for $i=0,\pm 1,\pm 2,\pm 3,\pm 4$ are also projectors, satisfying the orthogonality condition $P_{i}P_{j}=\delta_{ij}P_{i}$.
@@ 9 Massless tensors with arbitrary rank
Here $(P_{k})^{A}{}_{B}$ is defined as the sum of all possible terms $\pi_{i_{1}}\pi_{i_{2}}\cdots\pi_{i_{n}}$ such that $\sum_{m=1}^{n}i_{m}=k$ where we assign $\pi_{+}\equiv\pi_{1}$ and $\pi_{-}\equiv\pi_{-1}$. Note that the sequence of multipliers is significant. Since $\pi_{0}$, $\pi_{+}$ and $\pi_{-}$ are projectors it follows that all expressions $P_{k}$ are also projectors, satisfying the orthogonality condition $P_{k}P_{q}=\delta_{kq}P_{k}$ where $(k,q=0,\pm 1,\pm 2,\cdots,\pm n)$. The multiplication factor $k$ in (5) is consequence of combinatorics.
@@ 9.3 Dimension of subspaces $d_{k}$
The subspace projected by $\pi_{0}$ has dimension $2$ while those projected by $\pi_{+}$ and $\pi_{-}$ each have dimension $1$. Since $d_{k}=d_{-k}$ we obtain following expressions for the dimensions $d_{k}$.
@@ 10 Components with highest helicity
In previous sections, we demonstrated that our principle field equations successfully reproduce all known cases with lower helicities $\lambda=0,1,2$. We now turn to a comparison of our framework with established results for higher helicity fields $\lambda\geq 3$.
The general theory of free fields with arbitrary helicity $n$ was first developed by Fronsdal [7]. We will show that Fronsdal’s formulation emerges as a special case of our approach when restricted to the highest helicity component.
@@ 10.2 Highest helicity tensor $h_{A}(k)$ symmetric under space inversion in the frame of standard momentum
11 Relations with previous work
In this section, we employ our framework to systematically recover fundamental results from established studies of highest-helicity fields.
11.1 The highest-helicity fields in coordinate representation necessarily satisfy the Fierz-Pauli constraints
Let us go from standard momentum frame to arbitrary frame and then to coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$. Then we will consider field $h^{A}(x)$